A 96% solution of sulfuric acid is to be diluted with distilled water to make a 90% solution of sulfuric acid. How many litres of the orginal solution and how many litres of water are necessary to make 12.0L of the 90% solution?
Answer by ankor@dixie-net.com(22740)

A 96% solution of sulfuric acid is to be diluted with distilled water to make
a 90% solution of sulfuric acid. How many litres of the original solution
and how many litres of water are necessary to make 12.0L of the 90% solution?
:
Let x = amt of water
then
(12-x) = amt of 96% sulfuric acid
:
.96(12-x) = .90(12)
:
11.52 - .96x = 10.8
:
11.52 - 10.8 = .96x
:
.72 = .96x
x = (.72/.96)
x = .75 liters of water added to 11.25 liters of sulfuric acid
;
:
Check by finding the amt of sulfuric acid (amt should be unchanged)
.96(11.25) = .90(12)
10.8 = 10.8